OpenStudy (anonymous):

how to solve 4-(1/x)-(2/x^2)=0

myininaya (myininaya):

This is a quadratic if you let u=1/x try it.

myininaya (myininaya):

Can yu solve 4-u-2u^2=0? or -2u^2-u+4=0?

OpenStudy (anonymous):

im not sure where "u" comes into play

OpenStudy (joannablackwelder):

You don't have to solve it that way, though you can if you want. You could also multiply the equation by the LCD.

OpenStudy (joannablackwelder):

To clear the fraction.

myininaya (myininaya):

Either way.

myininaya (myininaya):

@vanne29 I just let u=1/x and seen that u^2=1/x and made substitutions.

myininaya (myininaya):

To get a quadratic.

myininaya (myininaya):

oops u^2=1/x^2*

OpenStudy (anonymous):

if i use the LCD method would i get 4x^2-1x-2=0?

myininaya (myininaya):

yep

OpenStudy (anonymous):

& then i use \[x=-b \pm\frac{ \sqrt{b ^{2}}-4ac }{ 2a }\] and plug in those number right?

myininaya (myininaya):

I think you mean \[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

myininaya (myininaya):

And if is that then yes

OpenStudy (anonymous):

the answer would be x=\[\pm \frac{ \sqrt{33} }{ 8 }\]

OpenStudy (anonymous):

\[1\pm \frac{ \sqrt{33} }{ 8}\]

myininaya (myininaya):

you mean \[x=\frac{1\pm \sqrt{33}}{8}\]?